OpenStudy (anonymous):

The face value of a single discount note is $12,000. The discount is 6 1/2% for 90 days. Use ordinary interest. Calculate the following: A. Amount of interest charged for each note. B. Amount borrower would receive. C. Amount Payee would receive at maturity. D. Effective rate (to the nearest tenth percent).

OpenStudy (anonymous):

try out this formula: ordinary interest= principal * rate* time , where: principal=$12000 rate=0.065 time= 90 days divided by 365 1/4 days= 1.5 mnths/12 mnths=1.5/12=0.125

OpenStudy (anonymous):

that wuz for A. to find how much the borrower receives you add the ordinary interest to the principal

OpenStudy (anonymous):

effective rate= A(1+r/n)^nt where; A=principal r=rate n=number of times in a year t=number of years